A sample of Ethane (C2H6) has a volume of 45.6 L at a pressure of 100.9 kPa and a temperature of 37.8˚C. What is the number of moles of the Sample? R=8.31 L kPa/K mol


1
Expert's answer
2021-04-07T11:28:04-0400

P = 100.9 kPa

T = 273+37.8 = 310.8K

V = 45.6 L

R = 8.31 LkPa/K mol

PV = nRT

100.9 kPa × 45.6 L = n × 8.31(L kPa/K mol) × 310.8

n = 1.78 mol